6/5,7/10,8/15,9/20
Write a rule for the nth term

Found 2 solutions by stanbon, scott8148:
Answer by stanbon(75887)

6/5,7/10,8/15,9/20
Write a rule for the nth term
-----
a(n) = (n+5)/(5n) where n = 1,2,3....
======================================
Cheers,
Stan H.

Answer by scott8148(6628)

an = (n + 5) / (5n)